Find two real numbers that have
a sum of 8 and a product of 2.

Answer by mananth(16946)

x+y=8
xy=2
x= 2/y
plug valu of x in the first equation
2/y + y =8
2+y^2 /y =8
2+y^2=8y
y^2-8y+2=0
..
quadratic formula
y1= (-(-8)+sqrt(64-8))/2a
y1=(8+sqrt56)/2
y1=7.74
y2=(8-sqrt 56)/2
y2=0.26
the numbers are 7.74 & 0.26